OpenStudy (anonymous):

For any real numbers x and y, if x^4 - y^4 = 0 then x=y or x=-y.

OpenStudy (anonymous):

true.

OpenStudy (anonymous):

It seems obvious, since proof by assume will give us an expected result, however, we need to prove that this is only true for (x,y)∈ℝ. So, proof by contrapositive?

OpenStudy (anonymous):

x^4 - y^4 = 0 => (x^2+y^2)(x^2-y^2) = 0 => x^2 = -y^2 OR (x+y)(x-y) = 0 Hmmmm....first root is imaginary stuff? Second part is clear. x = -y OR x = y.

OpenStudy (anonymous):

I think the question is asking how do we prove that ∀((x,y)∈ℝ)|x^4-y^4=0↔{x=y}∪{x=-y}

OpenStudy (anonymous):

Oh....yeah. if x and y are real, yes. Clearly.

OpenStudy (anonymous):

So the solutions by GT and and mad are partial solutions that would probably pass in a calculus class, but not in analytics.

OpenStudy (anonymous):

it is a set of questions preceded with "Prove these propositions:"

OpenStudy (anonymous):

this is an intro to proofs class.

OpenStudy (anonymous):

Probabilistic proof by assume contra. Solve for n and r such that substituting nx and ry in for x and y for {x,y}≠1?

OpenStudy (anonymous):

If the "then" statement is true, there cannot be non-unit values for scalar transforms of x and y such that n and r are real values.

OpenStudy (anonymous):

this one i am definitely not following.

OpenStudy (anonymous):

Alright, if we know the proposition given, THEN the x=y or x=-y. I'm not sure if that's a universal THEN, but if it is, we need to prove that there are no values x≠y or x≠-y such that this is also true.

OpenStudy (anonymous):

okay. but if we do it the way GT did it doesn't come out to a concrete result.

OpenStudy (anonymous):

If x^4 - y^4 = 0 then (x^2+y^2) (x^2-y^2) = 0. So that factors to (x^2+y^2) (x+y) (x-y) =0. Therefore, x^2=y^2 and x=y. (x+y) = 0 and (x-y)=0 so x=-y and x=y.

OpenStudy (turingtest):

how do you get from (x^2+y^2) (x+y) (x-y) =0 to x^2=y^2 ? aren't you supposed to be doing induction in class?

OpenStudy (anonymous):

i am just using what i know thus far. induction is still fairly new to me so i wouldn't know how to properly use it in this proof.

OpenStudy (turingtest):

ok I got it...

OpenStudy (turingtest):

\[x^4-y^4=0\iff x=\pm y\]first step: check for x=1

OpenStudy (anonymous):

Let x = 1 such that 1^4 - y^4 = 0. 1-y^4 = 0. -y^4=-1. ....

OpenStudy (turingtest):

try difference of squares...

OpenStudy (anonymous):

aright so then (1 - y^2) (1+y^2) = 0 so 1=y -1=y

OpenStudy (turingtest):

much better, though note that you could have done difference of squares twice\[(1-y^4)=(1+y^2)(1-y^2)=(1+y^2)(1+y)(1-y)=0\]since 1+y^2>0 we only have to solve the last two sets of parentheses for 0\[y=\pm1=\pm x\]that is a more thorough way to do it ;)

OpenStudy (turingtest):

so now we just assume that this is true for some random n\[n^4-y^4=0\]from here out out this statement is gospel the final step is to show that this implies that it is true for n+1 this can be done by plugging in n+1 for x, then using difference of squares and solving the resulting factored equation if you do that correctly you will be able to show that y=n+1, which means the statement is true for all n why don't you try that for a bit?

OpenStudy (turingtest):

sorry, the statement is\[n^4-y^4=0\iff y=\pm n\]

OpenStudy (anonymous):

you cant stop after showing y = +,- 1 = +, - x?

OpenStudy (turingtest):

no because that is just for one particular x (x=1) we want to show it for all x if that first statement was enough to prove it for all x I could prove that 1+x=2 for all x by just doing the first step! clearly that is not enough...

OpenStudy (turingtest):

what is needed is to 1)show this statement is true for some particular value 2)assume that it is true for some random value n (which we have already shown) 3) (this is the most important part) show that those two things /necessarily imply/ that this is true for n+1 that is sufficient to prove the statement

OpenStudy (anonymous):

can you help with the result? i'm stuck.

OpenStudy (turingtest):

checking x=n+1 we have\[((n+1)^4-y^4)=((n+1)^2+y^2)((n+1)^2-y^2)\]\[=((n+1)^2+y^2)((n+1)+y)((n+1)-y)=0\]again we have that the last two sets of parentheses are the only ones that could possibly be zero, so we have\[n+1-y=0\to y=(n+1)=x\]\[n+1+y=0\to y=-(n+1)=-x\]which I think proves our statement for at least all positive n I suppose we could have made x=0 (that is obviously true as well)

OpenStudy (turingtest):

hm... I'm not sure I see how to prove it for negative numbers, which seems to be required in the problem sorry, I'll have to think about that part

OpenStudy (anonymous):

this has to be done by induction?

OpenStudy (turingtest):

I was asking you if you were doing induction in class no, nothing has to be done that way, it's just a method it's probably the most common one for proving a statement like that about all numbers though

OpenStudy (anonymous):

could this be done just as a simple proof

OpenStudy (turingtest):

I'm not sure I know what that means... a proof requires what it requires ok here's a real cut-corner 'proof': you are trying to show that \[x^4-y^4=0\iff y=\pm x\] for all x and y if you think that just saying\[x^4-y^4=(x^2+y^2)(x+y)(x-y)=0\]the square of any number besides zero is positive, so x^2+y^2>0 therefor by the zero factor property either\[x+y=0\to x=-y\]or\[x-y=0\to x=y\]

OpenStudy (turingtest):

scratch the 'if you think that' part*